A certain ideal gas has a constant R = 38.9 ft-lf f/lb m-R with k = 1.4. a) If 3 lb m of this gas undergoes a reversible nonflow constant volume process from 120 psia,740 oF to 140 oF find P2, ∆U, ∆H, Q, Work. b) If the process in “a” had been steady flow find work and ∆S.


1
Expert's answer
2021-10-24T01:17:49-0400

Constant Volume Non Flow process


"P_2=P_1\times \frac{T_2}{T_1}= 120\times \frac{740}{140}= 634.28^0F"


"∆U = m\times C_v\times(T_2-T_1)= 3\times0.718 \times(740-140)=1292.4J =1.3kJ"


"∆H = \triangle U =1.3 kJ"


"q = \frac{Q}{m} =\frac {\triangle U}{m} =\frac{1292.4}{3}=430.8 J"

"W = \int p dV =0"